Two charges, Q and -Q, are held stationary at a distance of 4 m. Q= 5x10^-6 is given. a) What is the potential difference VB-VA? b) A point charge of mass 0.3 g and charge 2x10^-6 C, initially at rest, leaves point A. What is the velocity at B?


(The distance between Q and point A is 2 m. The distance between A and B are 1 m. The distance between B and -Q is 1 m.)


1
Expert's answer
2022-03-14T09:46:00-0400

a)

"\Delta \varphi=\frac{2kQ}r=22.5~kV,"

b)

"kqQ(\frac1{r_A}+\frac 1{r_B})=\frac{mv^2}2,"

"v=\sqrt{\frac{2kqQ(r_A+r_B)}{mr_Ar_B}}=90~\frac ms."